OpenStudy (gabylovesyou):

is this correct? Gino designed a garden for his rectangular backyard. The design he drew on paper is 6 inches wide. The width of his backyard is 30 feet. If his drawing is to scale, and his backyard is 40 feet long, how many inches long is his drawing? Write a proportion using words. Then, use the proportion to calculate the drawing length. Show your work. ---->(My answer) his scale is 6 inches to 30 feet = 1 inch for every 5 feet. if the back yard is 40 feet long, his drawing will be 40 feet (1 inch / 5 feet) = 8 inches long.

OpenStudy (campbell_st):

thats correct

jimthompson5910 (jim_thompson5910):

Yes you are correct, nice job You can also solve the following to get the same answer 30/6 = 40/x 30x = 6*40 30x = 240 x = 240/30 x = 8 Giving you the length of the backyard (on the drawing) in inches.

OpenStudy (gabylovesyou):

Thank you!! @campbell_st and @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (campbell_st):

we should be giving you a medal gaby nice work

OpenStudy (gabylovesyou):

Thank you!!!!! :D

jimthompson5910 (jim_thompson5910):

np